Matlab and Image Processing with MATLAB

1 Beginnings (Jan. 21/28; Feb. 4/11))

General references:

MATLAB Primer (Sixth Edition) Kermit Sigmon/Timothy A. Davis Chapman & Hall/CRC, 2002

Mastering MATLAB7 (A Comprehensive Tutorial and Reference) Duane Hanselman/Bruce Littlefield Pearson/Prentice Hall, 2005

Matlab Guide Desmond J. Higham/Nicholas J. Higham SIAM, 2000

Graphics and GUIs with MATLAB (Third Edition) Patrick Marchand/O. Thomas Holland Chapman & Hall/CRC, 2003

Training from the Mathworks: MATLAB Fundamentals and Programming Techniques

Matlab Programming for Engineers (2nd Edition) Stephen J. Chapman Brooks/Cole, 2002

Scientific Computing with MATLAB Alfio Quarteroni/Fausto Saleri Springer, 2003 MATLAB for Engineers Explained Fredrik Gustafsson/Niclas Bergman Springer, 2003

Essential MATLAB for Scientists and Engineers (2nd Edition) Brian D. Hahn Elsevier, 2003

Introduction to Matlab7 for Engineers William J. Palm III McGraw-Hill, 2005

All the pdf files indicated below are available at http://cda.psych.uiuc.edu/matlab pdf (a) Getting Started (getstart.pdf; handed out as hard copy in class)

1) Introduction

2) Matrices and Arrays

6) Desktoptools and Development Environment Webinar: Introduction to MATLAB 7 (Live broadcast was

Jan. 13, 2005) Matlab pdfs:

matlab env.pdf (desktoptools and development environment) math.pdf refbook.pdf refbook2.pdf refbook3.pdf

(b) Programming 4) Getting Started: Programming

Matlab pdfs:

matlab prog.pdf programming tips.pdf apiext.pdf apiref.pdf

(c) Graphics 3) Getting Started: Graphics

Matlab pdf:

graphg.pdf

(d) Guis 5) Getting Started: Creating Graphical User Interfaces

Matlab pdf:

buildgui.pdf

2 Statistics/Curve Fitting (Feb. 18/25; March 4)

References:

Computational Statistics Handbook with MATLAB Wendy L. Martinez/Angel R. Martinez Chapman & Hall/CRC, 2002

Exploratory Data Analysis with MATLAB Wendy L. Martinez/Angel R. Martinez Chapman & Hall/CRC 2005

Training from the Mathworks: Statistical Methods in MATLAB

(a) Webinar: Data Analysis with Curve Fitting and Statistics Toolboxes Matlab pdfs:

stats.pdf (Matlab Statistics Toolbox Manual) curvefit.pdf (Matlab Curve-Fitting Toolbox Manual)

(b) Webinar: Biomedical Statistics and Curve Fitting with MATLAB; New Statistics Tools in MATLAB and the Statistics Toolbox

(c) Webinar: MATLAB for Bioinformatics Matlab pdfs:

bioinfo ug.pdf gads tb.pdf (genetic algorithm)

3 Optimization (March 11/18)

References:

Applied Optimization with MATLAB Programming P. Venkataraman Wiley, 2002

Training from the Mathworks: MATLAB Based Optimization Techniques

(a) Webinar: Applied Optimization Using Matlab Matlab pdf:

optim tb.pdf (Matlab Optimization Toolbox Manual)

(b) Case study in changing a Fortran Subroutine to Matlab: The Dykstra-Karzmarz Fortran routine to MATLAB — (lin-

ear inequality constrained least-squares)

Case study in producing and calling Fortran dlls with MAT-LAB:

Fitting/Finding a unidimensional scaling for a proximity matrix

(c) Webinar: Genetic Algorithms and New Optimization Methods in MATLAB

4 Image Processing (April 1/8/15)

References:

Introduction to Digital Image Processing with MATLAB Alasdair McAndrew Thomson, 2004

Digital Image Processing Using Matlab Rafael C. Gonzalez/Richard E. Woods/Steven L. Eddins Pearson/Prentice Hall, 2004

Training from the Mathworks: MATLAB for Image Processing

(a) Webinar: Image Processing Using MATLAB Matlab pdf:

images tb.pdf (Matlab Image Processing Toolbox Manual)

(b) Webinar: Biomedical Image Processing with MATLAB

(c) Webinar: Geospatial Data Analysis Using MATLAB Matlab pdf:

map ug.pdf (Matlab Mapping Toolbox Manual)

5 Other Things (April 22/29)

(a) fMRI-type Packages SPM2 Lyngby Brainstorm EMEGS

(b) Various Related Toolboxes: Neural Net (nnet.pdf) Wavelet (wavelet ug.pdf) Spline (splines.pdf) Virtual Reality (vr.pdf) Symbolic (symbolic tb.pdf)